OpenStudy (anonymous):

.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? I give medals ^__^

OpenStudy (ajprincess):

Let the number of adults be a and the number of students be s. a+s=150 a=150-s 45a+28.5s=4770 45(150-s)+28.5=4770 solve for s and then substitute its value in a+s=150 to find a.